a body travels distance of 10m and 20m at constant velocity of 5m/s and 8m/s respectively, then what is its average velocity?...answer options are 6.5m/s , 8m/s, 6m/s, 20/3m/s..pls explain the answer...thanks in advance

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
Average velocity = 
= (10 + 20 )/[10/5 + 20/8]
= 30 /(2 + 2.5) = 30 /4.5
= 6.66 m/sec
Regards
Arun (askIITians forum expert)
Suryansh singh
one year ago
A body travels distance 10 m and 20 m at constant velocity of 5 m/s and 8 m/s respectively in same direction continuously, then average velocity of body is